Make New Entry,
Make Followup Entry
User name M Friend
Log entry time 07:30:17 on October 7,2009
Entry number 293610
keyword=Compton coincidence runs
We have 50uA beam, so we will try one more low current coincidence run. EDet moved to 1.0mm, 2 plane coincidences, prescales set to 4%, 0%, 100%.
**Run 20205 50uA coincidence run. Not stuck!!! Nice long run.
**Run 20206 100uA coincidence run. Rate not as high as I would expect.
Noticed that the edet motion ioc had rebooted itself, including the power being on now. Moved edet back to garage.